### v1

@@ 6.3 Experiment 3: Clustering Accuracy Test
Data: As input data, we used the same set of models used in experiment 1. In this experiment, we executed a mining task on the synthetic dataset adopted for experiment 1 and selected 33 out of the whole set of generated patterns, where each selected pattern can be easily traced back to one of the patterns presented in Figure 4 or one of its variations (e.g., Subkind patterns with a missing Subkind). We then manually clustered all the generated patterns by assigning a cluster label to them (e.g., we labeled the Subkind pattern and its variations as “cluster_0”). As the final segmentation, we produced six clusters, each one representing a pattern of Figure 4.
@@
Results: We depict the outcomes in Figure 6. For the dataset we considered, the clustering component we employ in the pipeline achieves 0.93 accuracy when the adopted similarity threshold is 0.6 or 0.61. The increase in accuracy in this is due to the creation of the same number of clusters, such that the different patterns have been distinguished and the variations for each of them not deemed as different patterns. The approach incorrectly predicted only two patterns in the best case,, as shown in the confusion matrix in the figure.
@@
Threats to validity: A primary concern regarding the validity of this experiment is that we conducted the test over a limited set of patterns by considering one clustering reference only. Users may vary in their way of classifying patterns, and it could be more challenging when the structures to be classified cannot be straightforwardly traced back to a reference pattern. Still, the test was enough to demonstrate how the clustering approach is reliable in distinguishing the structures that are generated and can be used to organize the data in a manner similar to what the user would adopt.
7 Utility Demonstration
@@
To illustrate the applicability and utility of our proposed approach, we conducted a simulation of a use case. The primary objective of this simulation was to showcase the effectiveness of our approach in analyzing the practical usage of a particular modeling language: OntoUML. Our specific goal was to demonstrate how the discovery pipeline can be used to extract valuable insights that can enhance our understanding of the language, facilitate modifications, and drive improvements. Within this scope, we have used an extended version of the OntoUML dataset, comprising 143 models. To ensure a substantial number of patterns for analysis, we conducted five comprehensive tests, adjusting parameters as needed to avoid generating an overwhelming number of structures. In a second step, we went through the whole set of generated patterns and we engaged in internal discussions with two experts who contributed to the design of the language.111111Note that we have deliberately separated this section from the experiments section. The reason is that in this case we did not properly design an experiment, but simulated a use case with two language designers, who also participated in the requirements definition and design of the approach we are offering. That’s why we talked about “demonstration” and “internal” discussion. The interaction with the experts occurred in two stages. First, we presented the generated patterns and asked them to identify the patterns for which it was worth having a more in-depth discussion. Second, after defining a subset of 40 patterns, we went through a discussion of each of them, focusing primarily on .i the unexpectedness of the mined structure, and .ii the possibility of reusing the structure somehow. This simulation allowed us to address a new research question:
@@
To what extent can the approach be used to discover new interesting structures? This research question relates to R1, R2 and R3, and explores the approach’s capability to identify compelling structures—potentially unexpected ones—in real-world scenarios featuring a diverse set of models with varying levels of complexity. If identified, these structures can offer valuable insights for designers, potentially informing their future design strategies.
@@
Figure 7 shows a structure with three Modes, each one of them characterizing a distinct bearer, where two Modes are subtypes of a parent Mode.121212Nodes with dashed line denoted as class_0 or class_1, i.e., without associated stereotype, are nodes that in the discovered patterns present relationships about which we have information related to the source node, but not the target node, or vice versa. This is because, as mentioned above, the graphs given as input to the mining algorithm have all relationships reified. So, for example, in Figure 7 class_1 was created because in the pattern we had a mode node associated with an edge source to a relation characterization.
The information related to the occurrence of the pattern (.b) reveals that the reference berares are represented by a Role, a Category, and a Kind, respectively. This does not yet present an actual problem; however, it may involve the generation of scenarios that are not intended. In fact, the pattern suggests checking whether role and kind are subtypes of the same concept. In fact, if this is not the case, the child modes might end up associated with two different bearers (e.g., class_0 and class_1 may classify different instances). This specific situation may be avoided by forcing the bearers of the child modes (e.g., class_0 and class_2) to be subclasses of the bearer of the main parent mode. In conclusion, the recurring structure we found does not present an incorrect practice, but it may trigger an evolution of the language. Originally, in fact, the designers of OntoUML did not dwell on the possibility of constructing taxonomies of modes. However, through the results of this mining activity, emerges how the de facto modelers use these taxonomies (the importance of which was also discussed in [40]). This uncovered structure, then, opens the door to introducing new possible language constraints to avoid unwanted configurations.
@@
That being considered, the OntoUML RoleMixin construct requires that there are at least two identity providers involved and then at least two different Role instances. Thus, in what sense does a RoleMixin participate in an Event? Looking at the occurrence of the pattern, the concept of RoleMixin is modeled as “client” and in this sense, we may have different instances of this class that are, for instance, persons or organizations. Considering this structure, to avoid possible mistakes, the cardinalities of the relationship between the RoleMixin concept and the Event concept must be made explicit. The recurrent structure we found indicates that the modeler left this information implicit. Consequently, the possible insight that can be collected by the language designer through the assessment of this structure is that the modeler should be guided in making explicit the connection between the identity providers and the given Event, possibly by providing the right cardinalities of the associations in question.
@@ 8 Discussion
RQ1: Can the proposed approach identify previously recognized interesting patterns? The answer to this question is affirmative. Experiment 1 demonstrated that already-known structures can be discovered accounting for all the constructs of which they are composed and without missing any occurrence. For the patterns we selected, we discovered all classes of patterns and their occurrences with 100% accuracy. The patterns we have chosen were also useful to demonstrate how the approach can account for the core constructs in the input language. The experiment also highlighted the importance of customization options is important in reducing the number of results, thus facilitating the identification process. The importance is enlarged when the methodology is applied to extensive real-world datasets, such as the one represented by the OntoUML catalog.
@@
RQ2: Which are the main parameters affecting the performance of the proposed approach? According to Table 3, the increase in pattern size and minimum support value improves the average performance, but decreases the number of patterns that can be found. Unlike when the goal is to search for more frequent structures, we can conclude from this that when it is necessary to find more information or possibly unexpected information, it is necessary to find the right trade-off between effectiveness and performance. On a dataset comprising OntoUML models, if we search too small and too infrequent patterns, the algorithm can produce even more than 10,000 outputs, often very similar to each other. Here, too, the role of customization steps can be essential, especially in reducing the information in the graphs to be sent as input to the mining algorithm and in eliminating redundant or unwanted outputs, thus also reducing the time of visualization and assessment.
@@
RQ3: Is the clustering step accurate in grouping patterns? Thanks to Experiment 3, we observe how our proposed clustering component provides a grouping criterion for patterns akin to manual user-performed ones. Our current configuration enables the differentiation of various structures by considering both the node count and the types of constructs involved. This capability positions it as a valuable asset during the results assessment phase. However, a point of discussion centers on the heterogeneity of the threshold values needed to achieve a satisfactory grouping, which, as we have found in other tests, can vary from case to case. We cannot assert a universally optimal threshold, such as adhering to a specific value like 5, for attaining the best grouping. Our trials have shown that optimal thresholds may range around 0.6 or 0.7 in some instances. One consistent observation, though, is that the optimal range for classification consistently falls within the mid-range of the scale. That is, thresholds between 0.1 and 0.3 or between 0.8 and 1.0 do not consistently achieve the most effective clustering. Considering the iterative nature of the discovery process, where users often experiment and fine-tune parameters to get the desired output in terms of size and type of structures, adjusting threshold values to achieve the desired clustering is also part of the iterative discovery process.
@@
RQ4: To what extent is the approach useful for discovering new interesting structures? From the tests we did for the utility demonstration, we observed that to find unexpected modeling practices, the importance value of frequency (or minimum support) decreases. In fact, if parameters are set to find frequent structures, it is more likely that we will find already known structures, especially with a language like OntoUML, where modelers operate following predefined guidelines and patterns. This mainly has a negative effect on the overall performance of the approach. In fact, as we saw in Experiment 2, the more we decrease the frequency threshold, the longer the mining process takes. Moreover, the approach has the potential to generate patterns of a very large order of magnitude. However, this did not prevent us from collecting useful and interesting information, such as that discussed in the previous section, which can trigger a range of analyzes and possibly interventions by the language designers that they had not thought of before. Currently, our emphasis is not on exploring the interesting types and unexpected structures that can be found, as it is not the main purpose of the paper. Our understanding of classifying these outputs is growing as we consider the analyzes and operations they might trigger.
@@
Structures triggering the definition of new patterns or anti-pattern. For instance, structures like the one represented in Figure 7 could be a good start for understanding and, possibly, devising new modeling strategies to be suggested to modelers. In the specific case of the example given, the structure of modes is a concrete case for understanding how to enable the designer to avoid generating unintended instantiations. For example, designers can concretely achieve this by exploiting the found structure and offering an ideal way of modeling modes taxonomies, or by introducing new constraints that force the creation of generalization links between the concepts characterized by the sub-modes and the concept characterized by the parent-mode.
@@ 9 Related Work
There is extensive literature [43, 44, 45, 29] on pattern discovery and its applications in a variety of domains, including software code [46], databases [47], educational processes [48], business processes [49], etc. However, the use of pattern discovery techniques is notably restricted in conceptual models. In this focused area of research, the closest work to what we propose is that of Skouradaki et al. [8], who designed a pattern mining algorithm for BPMN. Still, the goal of our contribution is not to provide a new mining algorithm. Our emphasis predominantly lies in the combination of a well-established FSM technique with graph manipulation techniques. Furthermore, a considerable amount of effort from our side concerns the definition of an interactive process where users can participate in the discovery activities, thus affecting the reliability of the final output. Last, we designed the approach with the scope of covering different conceptual modeling languages by keeping all the functions of the approach as language-independent.
@@
Two other approaches that are worth mentioning are the ones presented by Ardimento et al. [51] and the one that was introduced by Fumagalli et al. [52]. Both refer directly to UML class diagrams as the reference modeling language to be mined. However, the former is more aimed at mining modeling events that occur on Visual Paradigm,131313https://www.visual-paradigm.com/download/community.jsp i.e., the reference editing tool, and not properly frequent subgraph structures. The second, on the other hand, is much closer to our approach and is primarily focused on using frequent item set mining, instead of frequent subgraph mining, and also does not introduce the set of facilities we have offered here to support the discovery process.
To summarize, our approach distinguishes itself from related work through two key novel contributions. First, we empower users in their discovery process by offering a comprehensive set of interactive steps, providing guidance and assistance throughout the discovery journey. Second, we employ frequent subgraph mining, a robust established mining technique, for mining subgraphs in highly complex UML class diagrams represented by the models hosted in the OntoUML dataset. Our approach accounts for all vital elements of these diagrams, including cardinalities, generalization sets, multiple edge labels and class labels, and multi-directionality, thus ensuring the possibility of finding patterns containing all this information.

### v2

@@ 6.3 Experiment 3: Clustering Accuracy Test
The goal of this final experiment is to address RQ3 by testing the utility of the clustering component. It is important to note that the clustering component is not the primary focus of our contribution, so this is not the context to compare our implementation choices with existing alternatives (particularly regarding the embedding techniques and similarity computations used). Instead, the purpose of this test is mainly to assess how effectively this feature assists users in performing an output grouping task that would otherwise require manual effort.
6.3.1 Using the OntoUML Dataset
Data: As input data, we used the same set of models used in experiment 1. In this experiment, we executed a mining task on the synthetic dataset adopted for experiment 1 and selected 33 out of the whole set of generated patterns, where each selected pattern can be easily traced back to one of the patterns presented in Figure 7 or one of its variations (e.g., Subkind patterns with a missing Subkind). We then manually clustered all the generated patterns by assigning a cluster label to them (e.g., we labeled the Subkind pattern and its variations as cluster_0). As the final segmentation, we produced six clusters, each one representing a pattern of Figure 7.
@@ 6.3.1 Using the OntoUML Dataset
Results: We depict the outcomes in Figure 11. For the dataset we considered, the clustering component we employ in the pipeline achieves 0.93 accuracy when the adopted similarity threshold is 0.6 or 0.61. The increase in accuracy in this is due to the creation of the same number of clusters, such that the different patterns have been distinguished and the variations for each of them not deemed as different patterns. The approach incorrectly predicted only two patterns in the best case, as shown in the confusion matrix in the figure.
The confusion matrix (right) was reordered using the Hungarian algorithm,181818https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.linear_sum_assignment.html to optimize alignment between true and predicted clusters, ensuring a clearer visualization of results and accounting for the fact that cluster labels are assigned arbitrarily by the clustering algorithm (e.g., predicted cluster_6 could represent ground truth cluster_1). Rows represent true clusters (ground truth labels), columns represent predicted clusters and diagonal cells indicate correctly predicted samples, while off-diagonal cells highlight misclassification. The confusion matrix reports the results at the optimal thresholds (0.60–0.61).
6.3.2 Using the ArchiMate Dataset
Data: For this experiment, we used the same dataset of ArchiMate models as in experiment 1 (see Section 6.1.2) and we selected 20 patterns from the mining output of the neutral trial in experiment 1. Each selected pattern corresponds to one of the patterns depicted in Figure 9 or a variation (e.g., missing an element or relationship). The chosen patterns were manually clustered into groups to serve as ground truth in our evaluation (e.g., Chatty Service patterns are assigned the label cluster_0), resulting in six total clusters, each representing a pattern of Figure 9.
Setup: We evaluated the clustering component by applying multiple similarity thresholds (e.g., 0.1, 0.2, …, 0.9) and comparing the clustering outputs with the ground truth. Initially, to calculate clustering accuracy, we used the same pairwise comparison method as in the OntoUML experiment (see Section 6.3.1). For each pair of patterns, the predicted clustering result was compared with the ground truth to determine if the clustering algorithm accurately grouped or separated them. To refine the evaluation, we further explored thresholds within the interval where the accuracy peaked (e.g., 0.65–0.7) by testing intermediate values (e.g., 0.66, 0.67).
Results: The results of this experiment are shown in Figure 12. The bar plot (left) shows pair-wise accuracy (from Section 6.3.1) across various clustering thresholds. The highest accuracy, 0.96, was achieved within the range of 0.65–0.7. Beyond this range, accuracy slightly decreased due to over-clustering or merging distinct clusters. The confusion matrix (right) was reordered using the Hungarian algorithm to optimize alignment between true and predicted clusters, as in the OntoUML scenario. The results reported concern the optimal threshold between 0.65 and 0.7.
6.3.3 Threats to Validity
@@
A primary concern regarding the validity of this experiment is that we conducted the test over a limited set of patterns by considering one clustering approach only.191919Note that for the current implementation, we adopted a straightforward approach where all vectors must be of the same size, with the same dimensions. Each pattern is represented as a record in a matrix with features such as “number of nodes”, “kind”, “number of relations”, “general”, and “specific”, etc. Each feature can hold different values for each pattern (for instance, if a pattern has three nodes classified as “kind”, it will have “number of nodes” = 3 and “kind” = 3, while if it lacks nodes for generalizations, it will report 0 for both “specific” and “general”). Users may vary in their way of classifying patterns, and it could be more challenging when the structures to be classified cannot be straightforwardly traced back to a reference pattern. Still, the test was enough to demonstrate how the clustering approach is reliable in distinguishing the structures that are generated and can be used to organize the data in a manner similar to what the user would adopt.
7 Application Example
@@
To illustrate the applicability and utility of our proposed approach, we conducted a simulation of a use case. The primary objective of this simulation was to showcase the effectiveness of our approach in analyzing the practical usage of a particular modeling language: OntoUML. Our specific goal was to demonstrate how the discovery pipeline can be used to extract valuable insights that can enhance our understanding of the language, facilitate modifications, and drive improvements. Within this scope, we have used an extended version of the OntoUML dataset, comprising 143 models.202020More information about this ever-growing dataset, with statistics about models, can be found in [15] and https://github.com/OntoUML/ontouml-models. To ensure a substantial number of patterns for analysis, we conducted five comprehensive tests, adjusting parameters to avoid generating an overwhelming number of structures.212121Note that the mining process should be viewed as an iterative procedure, concluding once interesting results are obtained, so it becomes difficult to envision an optimal set of parameter a priori. Testing the current approach to establish optimal mining parameters and exploit the pipeline to generate a robust dataset of patterns could be a valuable objective for future research. In a second step, we went through the whole set of generated patterns and we engaged in internal discussions with two experts who contributed to the design of the language.222222Note that we have deliberately separated this section from the experiments section. The reason is that in this case we did not properly design an experiment, but simulated a use case with two language designers, who also participated in the requirements definition and design of the approach we are offering. That’s why we talked about “demonstration” and “internal” discussion. The interaction with the experts occurred in two stages. First, we presented the generated patterns and asked them to identify the patterns for which it was worth having a more in-depth discussion. Second, after defining a subset of 40 patterns, we went through a discussion of each of them,232323The examples in this section can be viewed as a summary of discussions on the most relevant patterns. We added this clarification as a note at the location indicated by the reviewer. focusing primarily on .i the unexpectedness of the mined structure, and .ii the possibility of reusing the structure somehow. This simulation allowed us to address a new research question:
@@
To what extent can the approach be used to discover new interesting structures? This research question relates to R1, R2 and R3, and explores the approach’s capability to identify compelling structures—potentially unexpected ones—in real-world scenarios featuring a diverse set of models with varying levels of complexity. If identified, these structures can offer valuable insights for language engineers, potentially informing their future design strategies.
@@
Figure 13 shows a structure with three Modes, each one of them characterizing a distinct bearer, where two Modes are subtypes of a parent Mode.242424Nodes with dashed line denoted as class_0 or class_1, i.e., without associated stereotype, are nodes that in the discovered patterns present relationships about which we have information related to the source node, but not the target node, or vice versa. This is because, as mentioned above, the graphs given as input to the mining algorithm have all relationships reified. So, for example, in Figure 13 class_1 was created because in the pattern we had a mode node associated with an edge source to a relation characterization.
The information related to the occurrence of the pattern (.b) reveals that the reference bearers are represented by a Role, a Category, and a Kind, respectively. This does not yet present an actual problem; however, it may involve the generation of scenarios that are not intended. In fact, the pattern suggests checking whether role and kind are subtypes of the same concept. In fact, if this is not the case, the child modes might end up associated with two different bearers (e.g., class_0 and class_1 may classify different instances). This specific situation may be avoided by forcing the bearers of the child modes (e.g., class_0 and class_2) to be subclasses of the bearer of the main parent mode. In conclusion, the recurring structure we found does not present an incorrect practice, but it may trigger an evolution of the language. Originally, in fact, the designers of OntoUML did not dwell on the possibility of constructing taxonomies of modes. However, through the results of this mining activity, emerges how the de facto modelers use these taxonomies (the importance of which was also discussed in [47]). This uncovered structure, then, opens the door to introducing new possible language constraints to avoid unwanted configurations.
@@
That being considered, the OntoUML RoleMixin construct requires that there are at least two identity providers involved and then at least two different Role instances. Thus, in what sense does a RoleMixin participate in an Event? Looking at the occurrence of the pattern, the concept of RoleMixin is modeled as “client” and in this sense, we may have different instances of this class that are, for instance, persons or organizations. Considering this structure, to avoid possible mistakes, the cardinalities of the relationship between the RoleMixin concept and the Event concept must be made explicit. The recurrent structure we found indicates that the modeler left this information implicit. Consequently, the possible insight that can be collected by the language engineer through the assessment of this structure is that the modeler should be guided in making explicit the connection between the identity providers and the given Event, possibly by providing the correct cardinalities of the associations in question.
@@ 8 Discussion
RQ1: Can the proposed approach generate structures encoding previously recognized interesting patterns? The answer to this question is affirmative. Experiment 1 demonstrated that already-known structures can be discovered accounting for all the constructs of which they are composed and without missing any occurrence. For the patterns we selected, we successfully discovered all classes of patterns and their occurrences. The patterns we have chosen were also useful to demonstrate how the approach can account for the core constructs in the input language. The experiment also highlighted the importance of customization options is important in reducing the number of results, thus facilitating the identification process. The importance is enlarged when the methodology is applied to extensive real-world datasets, such as the one represented by the OntoUML catalog.
@@
RQ2: Which are the main parameters affecting the performance of the proposed approach? According to Table 5 and Table 6, the increase in pattern size and minimum support value improves the average performance but decreases the number of patterns that can be found. Unlike when the goal is to search for more frequent structures, we can conclude from this that when it is necessary to find more information or possibly unexpected information, it is necessary to find the right trade-off between effectiveness and performance. On a dataset comprising OntoUML or ArchiMate models, if we search too small and too infrequent patterns, the algorithm can produce even more than 10.000 outputs, often very similar to each other. Here, too, the role of customization steps can be essential, especially in reducing the information in the graphs to be sent as input to the mining algorithm and in eliminating redundant or unwanted outputs, thus also reducing the time of visualization and assessment.
@@
RQ3: Is the clustering step accurate in grouping structures? Thanks to Experiment 3, we observe how our proposed clustering component provides a grouping criterion for patterns akin to manual user-performed ones. Our current configuration enables the differentiation of various structures by considering both the node count and the types of constructs involved. This capability positions it as a valuable asset during the results assessment phase. However, a point of discussion centers on the heterogeneity of the threshold values needed to achieve a satisfactory grouping, which, as we have found in other tests, can vary from case to case. We cannot assert a universally optimal threshold, such as adhering to a specific value like 5, for attaining the best grouping. Our trials have shown that optimal thresholds may range around 0.6 or 0.7 in some instances. One consistent observation, though, is that the optimal range for classification consistently falls within the mid-range of the scale. That is, thresholds between 0.1 and 0.3 or between 0.8 and 1.0 do not consistently achieve the most effective clustering. Considering the iterative nature of the discovery process, where users often experiment and fine-tune parameters to get the desired output in terms of size and type of structures, adjusting threshold values to achieve the desired clustering is also part of the iterative discovery process.
@@
RQ4: To what extent is the approach useful for discovering new interesting structures? From the tests we did for the utility demonstration, we observed that to find unexpected modeling practices, the importance value of frequency (or minimum support) decreases. In fact, if parameters are set to find frequent structures, it is more likely that we will find already known structures, especially with a language like OntoUML, where modelers operate following predefined guidelines and patterns. This mainly has a negative effect on the overall performance of the approach. In fact, as we saw in Experiment 2, the more we decrease the frequency threshold, the longer the mining process takes. Moreover, the approach has the potential to generate patterns of a very large order of magnitude. However, this did not prevent us from collecting useful and interesting information, such as that discussed in the previous section, which can trigger a range of analyses and possibly interventions by the language engineers that they had not thought of before. Currently, our emphasis is not on exploring the interesting types and unexpected structures that can be found, as it is not the main purpose of the paper. Our understanding of classifying these outputs is growing as we consider the analyzes and operations they might trigger.
@@
Structures triggering the definition of new patterns or anti-pattern. For instance, structures like the one represented in Figure 13 could be a good start for understanding and, possibly, devising new modeling strategies to be suggested to modelers. In the specific case of the example given, the structure of modes is a concrete case for understanding how to enable the engineers to avoid generating unintended instantiations. For example, engineers can concretely achieve this by exploiting the found structure and offering an ideal way of modeling modes taxonomies, or by introducing new constraints that force the creation of generalization links between the concepts characterized by the sub-modes and the concept characterized by the parent-mode.
@@ 9 Related Work
There is extensive literature [50, 51, 52, 32] on pattern discovery and its applications in a variety of domains, including software code [53], databases [54], educational processes [55], business processes [56], model-driven engineering (MDE) [57], EMF metamodels [58], etc. However, to the best of our knowledge, there is a significant potential for research into automatic applications that support pattern discovery in conceptual models.
In this focused area of research, the closest work to what we propose is that of Skouradaki et al. [12], who designed a pattern mining algorithm for BPMN. Still, the goal of our contribution is not to provide a new mining algorithm. Our emphasis predominantly lies in the combination of a well-established FSM technique with graph manipulation techniques. Furthermore, a considerable amount of effort from our side concerns the definition of an interactive process where users can participate in the discovery activities, thus affecting the reliability of the final output. Last, we designed the approach with the scope of covering different conceptual modeling languages by keeping all the functions of the approach as language-independent.
@@
Two other approaches that are worth mentioning are the ones presented by Ardimento et al. [61] and the one that was introduced by Fumagalli et al. [10]. Both refer directly to UML class diagrams as the reference modeling language to be mined. However, the former is more aimed at mining modeling events that occur on Visual Paradigm,252525https://www.visual-paradigm.com/download/community.jsp i.e., the reference editing tool, and not properly frequent subgraph structures. The second, on the other hand, is much closer to our approach and is primarily focused on using frequent item set mining, instead of frequent subgraph mining, and also does not introduce the set of facilities we have offered here to support the discovery process.
Along similar lines, particularly within the context of MDE and, more broadly, software engineering, other solutions exist that address problems similar to ours.
For instance, the work in [62] aims to enhance software maintenance by identifying design motifs, which are solutions to recurring design challenges. The authors employ two pattern-matching algorithms adapted from bioinformatics—automata simulation and bit-vector processing—to detect exact and approximate occurrences of design motifs in object-oriented code. The findings from this research underscore the potential of bioinformatics-inspired graph-matching methods to facilitate design recovery, streamline the analysis of complex software architectures, and uncover embedded design patterns.
Likewise, the study in [63] tackles the issues of recognizing design patterns in object-oriented programs. The proposed approach utilizes design motifs to identify pattern instances and filter out irrelevant matches. This strategy effectively narrows the search space, aiding the pattern discovery process.
To summarize, our approach distinguishes itself from related work through two key novel contributions. First, we empower users in their discovery process by offering a comprehensive set of interactive steps, providing guidance and assistance throughout the discovery journey. Second, we employ frequent subgraph mining, a robust established mining technique, for mining subgraphs in diagrams representing models in different languages. Our approach also accounts for all vital elements of these diagrams, including cardinalities, multiple edge labels and class labels, and multi-directionality, thus ensuring the possibility of finding patterns containing all this information.
As a final remark, the approach we propose employs various techniques across each module in the pipeline. For example, we use a technique for mining frequent structures, a technique for matching to trace domain information, a technique for transforming mining algorithm outputs into vectors, and a technique for clustering these vectors. Each of these techniques could merit its own discussion of related work, spanning fields such as “metamodel clone detection”, “linguistic corpus analysis”, “efficient pattern similarity computation and clustering”, and “plagiarism detection”.
However, we have focused our related work review on approaches that address our specific problem: the automatic support for identifying recurrent structure heuristics. Therefore, where relevant, we have included information on complementary techniques related to individual pipeline steps within the sections dedicated to implementation choices.